# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ III.6 Conversion between conservative and primitive variables
We need to convert from the conserved variables $(D,\tau,S_{i})$ to the primitive variables $(\rho,\epsilon,P,Wvi^{i})$ at every Runge-Kutta substep. This needs to happen because we need the primitive variables for computing flux vectors from Eq. (16), for setting source terms from Eq. (17), and the atmosphere treatment from Sec. III.5.1. Computing the conserved variables from the primitive ones is straightforward through Eqs. (11-13), but the reverse operation involves a root finder since the equations cannot be inverted analytically. The basic principle of conversion between primitive and conservative variables for the simulations in this paper remains similar to Sec. 4.3.1 of [76], with some small adjustments. We still use the same procedure described in [76] to solve for the root value of
@@
However, unlike in [76], first, our current procedure takes into account whether the point we are solving for is in the atmosphere region or not. If at any point we find $D<\rho_{\mathrm{atmo,cut}}$, we simply set $\rho_{0}=D$ and $\epsilon=P=Wv^{i}=0$, without going through the actual root solving procedure. This is different from the procedure in [76] as that method would perform the conversion everywhere through the root solver. By avoiding root solving in these low density regions, which is mostly the regions that result in unphysical states as mentioned before, we avoid the associated numerical complications that might cause the root solver to fail. Additionally, even when we do perform the actual root solving, we take certain precautions against possible complications. First, as root finding can fail due to round off error near the upper and lower limits of the interval in which the root must lie according to [116], we extend the maximum of the range to $1.1(Wv)_{\mathrm{max}}$, with $(Wv)_{\mathrm{max}}$ being the upper limit. Secondly, we set a floor value for $\epsilon_{\mathrm{floor}}$ such that if $\epsilon$ falls below this through the conversion process, we simply set $\epsilon=\epsilon_{\mathrm{floor}}$. The particular value for $\epsilon_{\mathrm{floor}}$ used is stated for the simulations in Sec. IV. The last point where we deviate from the prescription in [76] is that we do not perform any kind of rescaling if $Z>Wv$, where $Z=\displaystyle\sqrt{S^{i}S_{i}}/(Wh\rho_{0})$. Just like for the limiters, we also collect information about any trouble when converting from conserved to primitive variables at each point in an element during an evolution substep. The reasoning behind this is that in an element that has smooth and well-behaved matter fields, the conversion procedure should be carried out seamlessly. However, any kind of problem or difficulty during the conversion process will indicate that the matter fields are not well-behaved, and hence the element is marked as troubled. This information is then used in our trouble criteria list as mentioned in Sec. III.7.
@@ III.7 Hybrid DG+FV/FD evolution scheme
We want to use DG in the elements that are not troubled. In the troubled elements, we want to use FV to evolve the matter fields and FD to evolve the gravity fields. The concept of “trouble” is defined mainly based on the smoothness of the matter fields in an element. If the matter fields in an element are not smooth, the element is deemed “troubled”. Additionally, an element is also considered “troubled” if unphysicallities occur in the matter fields in the element either in low density regions or during conversion from conserved to primitive hydrodynamic variables. For determining trouble in an element, we use the following prescription:
i. We perform a full time evolution step from $t$ to $t+\Delta t$ through the RK time evolution scheme described in Sec. III.3 and at each RK substep, we collect information about troubles during the substeps. These include non-physical behavior that triggers one of the positivity limiters discussed in Sec. III.5.2, and problems when converting between conserved and primitive variables as described in Sec. III.6.
@@
    Attempt a conversion from the conserved to primitive variables for the upcoming time step in the element. If the attempt fails, consider element as troubled.
@@
    Look at the expansion coefficients of the fields in the basis polynomials and decide based on Persson method [117] in an element and demand
@@
    Hence, we should consider only the $N_{l,\mathrm{uf}}$ number of coefficients that remain unfiltered for each direction, as these are the ones that truly represent the smoothness of the fields in an element. In order to determine which coefficients have not been filtered, we only count those where the filter factor is above some factor $f_{\mathrm{unfiltered}}$, i.e., we only include $c_{l_{1}l_{2}l_{3}}$ if $c_{l_{1}l_{2}l_{3}}\geq f_{\mathrm{unfiltered}}c_{l_{1}l_{2}l_{3}}^{\mathrm{unfiltered}}$. For the choice of filtering in the simulations presented in Sec. IV, we find that the number of unfiltered coefficients $N_{l,\mathrm{uf}}$ is related to the number of total coefficients $N_{l}$ in each direction as $N_{l,\mathrm{uf}}=N_{l}-1$. Note that the coefficients are always computed using the DG grid points. This is straightforward in an element using the DG scheme, as we have the values of the variable we are checking already at the DG grid points. However, if we want to check the criterion in a FV/FD element, we first interpolate onto the DG grid points using a twelfth order Lagrange interpolation scheme, and then compute the expansion coefficients on this reconstructed DG grid. Lastly, $\alpha_{N}^{\mathrm{evo}}$ is a relaxation parameter. For our simulations, we set $\alpha_{N}^{\mathrm{evo}}=8$ in the DG elements and $\alpha_{N}^{\mathrm{evo}}=9$ in the FV/FD elements, except for the blast wave propagation tests in Sec. IV.2. We relax our parameters in Sec. IV.2, by using $4$ for DG elements and $5$ for FV/FD elements. The higher values in FV/FD elements is because we want to be more cautious when switching from FV/FD to DG as DG will not work well in a troubled element. From our experience, this Persson trouble indicator plays an especially crucial role in the accuracy of the simulation.
These criteria are checked in order, and if any one of them indicate trouble at any point in an element, the element is marked as troubled by the TEI and no further check is performed.
iii. If an element is marked troubled by the TEI, then if we are using DG, we switch to FV/FD and redo the full RK step. Conversely, if we are using FV/FD then we keep doing FV/FD for the next RK step. On the other hand, if the element is not marked troubled, then if we are using DG, we try doing DG for the next step. However, if we are using FV/FD, then we switch to DG only after eleven consecutive time steps with no trouble to be safe. In the DG elements, we use half the number of points as in the FV/FD elements, as DG can generate higher accuracy results at low computational cost.
@@ IV Numerical results
Table 1: Specifics of the different single neutron stars we simulate. The first column states the name and the abbreviation we assign. The second column states the initial central density. The third and fourth column state the rest mass and the initial star radius (in Schwarzschild coordinates). The fifth column states the strength of the initial perturbation (either in pressure $P$ or density $\rho_{0}$) and the sixth column states the speed imparted to the center of mass (CoM). The last column states the choices of evolution scheme we use to simulate the case. In each case, use the ideal EoS of Eq. (25) with $\kappa=100$ and $n=1$.
Name $\rho^{C}_{0}\cdot 10^{3}M_{\odot}^{2}$ $m_{0}/M_{\odot}$ $R_{\mathrm{Sch}}/M_{\odot}$ $\lambda$ $v_{\mathrm{CoM}}$ Scheme(s)
@@
In this paper we present simulations that we have performed to ensure the robustness and stability of our new methods. We start with a scalar wave test, as this helps us test our new DG+FV/FD scheme purely by itself, without the numerical complexities arising from matter fields in neutron star systems. Having checked proper convergence for this simple case, we move on to evolving initial data with special relativistic Reimann discontinuities in the initial data to test hydrodynamic shock propagation. Next, we simulate single neutron stars. We simulate four different cases of single neutron star initial data that are outlined in Table 1, while the resolutions used for each of the initial setup are described in Table 2.
@@
Table 2: Initial resolution and number and type of elements across the star for different single neutron star cases we simulate. First column is name of case as per Table 1 and the second column is the corresponding section where the case is discussed. $L$ in the third column states the dimension of the central cube. Fourth column states the evolution scheme or schemes used. Fifth and sixth column states the number of points in each direction in DG and FV/FD elements respectively, as applicable. Seventh column states center velocity $v_{\mathrm{CoM}}$. Eighth column specifies the highest level of h-refinement in the refinement region centered on the star. Last column states the initial total number of elements across the star, while specifying how many are of each type. When counting these elements, we also count those which partially contain the star at its surface. In the DG+FV/FD cases, the FV/FD elements are on the star surface, while DG elements are in the center. The elements for the $v_{\mathrm{CoM}}=0.5$ case are counted parallel ($\parallel$) to $\vec{v}_{\mathrm{CoM}}$ and perpendicular ($\perp$) to $\vec{v}_{\mathrm{CoM}}$ direction.
Name Sec. $L$ Scheme $n$(DG) $n$(FV/FD) $v_{\mathrm{CoM}}$ $l$(max) Star elements
(S) IV.3.1 $32$ DG $5$ 0 $4$ $10$ DG
$5$ $18$ DG
$6$ $34$ DG
IV.3.2 $52$ DG+FV/FD $8$ $16$ 0 $5$ $8$ DG $+$ $4$ FV/FD
(P) IV.4 $32$ DG $5$ 0 $4$ $10$ DG
$5$ $18$ DG
$6$ $34$ DG
(U) IV.5.1 $16$ FV/FD - $12$,$18$,$28$,$42$ 0 $3$ $5$ FV/FD
IV.5.2 $16$ FV/FD - $16$ 0 $3$ $5$ FV/FD
$4$ $9$ FV/FD
$5$ $17$ FV/FD
IV.5.3 $16$ DG+ FV/FD $8$ $16$ 0 $3$ $5$ FV/FD
$4$ $9$ FV/FD
(B) IV.6 $52$ DG+FV/FD $8$ $16$ $0.5$ $5$ $5$ DG $+$ $2$ FV/FD $\parallel\vec{v}_{\mathrm{CoM}}$,
$5$ DG $+$ $4$ FV/FD $\perp\vec{v}_{\mathrm{CoM}}$
$0.1$ $4$ $1$ DG $+$ $4$ FV/FD
$5$ $5$ DG $+$ $4$ FV/FD
@@ IV.5.3 $16$ DG+ FV/FD $8$ $16$ 0 $3$ $5$ FV/FD
Each of these cases are discussed in detail in the sections below. Out of these, (S) tests for stability and robustness, while the rest test the ability of the program to handle various types of dynamic scenarios. In order to study the behavior of the numerical results with increasing resolution, we use both p-refinement and h-refinement. P-refining an element simply involves increasing the number of points in the element in each direction as necessary. In case of h-refinement of an element, we take the element and split it in two in each direction. This results in eight new elements replacing the old element. Each of these eight new elements contain the same arrangement of points as the parent element. For DG elements, p-refinement increases the basis polynomial order, which would cause an exponential dropoff in the error in the results in the case of smooth fields, but can also prove problematic when non-smooth fields are present in an element, as these can give rise to sharp noisy features in the higher basis modes. In this case, for DG elements, it is more convenient to increase resolution using h-refinemet instead. For the case of pure FV/FD evolutions on the other hand, we can employ both p-refinement and h-refinement as necessary.
@@
To the best of our knowledge, this is the first time DG methods have been used for boosted neutron stars (B) and for unstable neutron stars (U) (that migrate to a stable state) in fully 3D simulations. While [87, 65, 97] evolve single TOV stars with a hybrid DG-FD scheme, they study the initially static configuration and a rotating configuration. Although simulations of the initially unstable configuration are performed with DG methods in [83], a 1D spherically symmetric implementation is used. This case setup is also studied in [114], where simulations are performed using spherically symmetric and also axisymmetric cartoon methods [118, 119]. As such, none of the prior simulations are in full 3D. Additionally, none of them evolve boosted neutron stars.
@@ IV.1 Tests with a smooth scalar field
In order to keep MPI communication at a minimum, our particular FV method does not use any extra ghost points (unlike more traditional finite volume codes). Communication has to happen only for points on the element surfaces as in the DG method. This is crucial for good strong scaling. On the other hand, our FV method becomes less accurate near element boundaries, because we need to drop to a lower order FV scheme there, due to the lack of ghost points. We have tested the method by evolving a smooth scalar field, using the same conservative evolution equations as described in Sec. 4.1 of [76]. We choose the mesh to be a 2-dimensional square that is equally divided into nine smaller square patches. In the middle patch we use FV, while the surrounding eight patches use DG. We then uniformly h-refine each patch, so that each of the nine patches is covered by an equal number of elements. In each element we use only 4 points (in each of the two directions), which is a challenging test for our FV method that works best if the number of points in the interior is larger than the number of element boundary points, because our method falls back to lower accuracy at element boundaries as explained in Sec. III.2.
@@
In Fig. 3 we show how the numerical error converges to zero with resolution. Our method is consistent with second order convergence. While second order convergence may not sound too impressive, we note that in shocked regions we cannot expect more than first order convergence in any case. Thus, our method is adequate when used in elements with shocks. Our plan is then to use this FV method only in the elements where the matter is not smooth, while we plan to use the DG method everywhere else. In this way we should get the benefits of both FV and DG as suggested in [87, 65].

### v2

@@ III.6 Conversion between conservative and primitive variables
We need to convert from the conserved variables $(D,\tau,S_{i})$ to the primitive variables $(\rho,\epsilon,P,Wvi^{i})$ at every Runge-Kutta substep. This needs to happen because we need the primitive variables for computing flux vectors from Eq. (16), for setting source terms from Eq. (17), and the atmosphere treatment from Sec. III.5.1. Computing the conserved variables from the primitive ones is straightforward through Eqs. (11)-(13), but the reverse operation involves a root finder since the equations cannot be inverted analytically. The basic principle of conversion between primitive and conservative variables for the simulations in this paper remains similar to Sec. 4.3.1 of [76], with some small adjustments. We still use the same procedure described in [76] to solve for the root value of
@@
However, unlike in [76], first, our current procedure takes into account whether the point we are solving for is in the atmosphere region or not. If at any point we find $D<\rho_{\mathrm{atmo,cut}}$, we simply set $\rho_{0}=D$ and $\epsilon=P=Wv^{i}=0$, without going through the actual root solving procedure. This is different from the procedure in [76] as that method would perform the conversion everywhere through the root solver. By avoiding root solving in these low-density regions, which is mostly the regions that result in unphysical states as mentioned before, we avoid the associated numerical complications that might cause the root solver to fail. Additionally, even when we do perform the actual root solving, we take certain precautions against possible complications. First, as root finding can fail due to round off error near the upper and lower limits of the interval in which the root must lie according to [117], we extend the maximum of the range to $1.1(Wv)_{\mathrm{max}}$, with $(Wv)_{\mathrm{max}}$ being the upper limit. Secondly, we set a floor value for $\epsilon_{\mathrm{floor}}$ such that if $\epsilon$ falls below this through the conversion process, we simply set $\epsilon=\epsilon_{\mathrm{floor}}$. The particular value for $\epsilon_{\mathrm{floor}}$ used is stated for the simulations in Sec. IV. The last point where we deviate from the prescription in [76] is that we do not perform any kind of rescaling if $Z>Wv$, where $Z=\displaystyle\sqrt{S^{i}S_{i}}/(Wh\rho_{0})$. Just like for the limiters, we also collect information about any trouble when converting from conserved to primitive variables at each point in an element during an evolution substep. The reasoning behind this is that in an element that has smooth and well-behaved matter fields, the conversion procedure should be carried out seamlessly. However, any kind of problem or difficulty during the conversion process will indicate that the matter fields are not well behaved, and hence the element is marked as troubled. This information is then used in our trouble criteria list as mentioned in Sec. III.7.
@@ III.7 Hybrid DG+FV/FD evolution scheme
We want to use DG in the elements that are not troubled. In the troubled elements, we want to use FV to evolve the matter fields and FD to evolve the gravity fields. The concept of “trouble” is defined mainly based on the smoothness of the matter fields in an element. If the matter fields in an element are not smooth, the element is deemed “troubled”. Additionally, an element is also considered “troubled” if unphysicalities occur in the matter fields in the element either in low-density regions or during conversion from conserved to primitive hydrodynamic variables. For determining trouble in an element, we use the following prescription:
i. We perform a full-time evolution step from $t$ to $t+\Delta t$ through the RK time evolution scheme described in Sec. III.3 and at each RK substep, we collect information about troubles during the substeps. These include nonphysical behavior that triggers one of the positivity limiters discussed in Sec. III.5.2, and problems when converting between conserved and primitive variables as described in Sec. III.6.
@@
    Attempt a conversion from the conserved to primitive variables for the upcoming time step in the element. If the attempt fails, consider the element as troubled.
@@
    Look at the expansion coefficients of the fields in the basis polynomials and decide based on Persson method [118] in an element by demanding
@@
    Hence, we should consider only the $N_{l,\mathrm{uf}}$ number of coefficients that remain unfiltered for each direction, as these are the ones that truly represent the smoothness of the fields in an element. In order to determine which coefficients have not been filtered, we only count those where the filter factor is above some factor $f_{\mathrm{unfiltered}}$, i.e., we only include $c_{l_{1}l_{2}l_{3}}$ if $c_{l_{1}l_{2}l_{3}}\geq f_{\mathrm{unfiltered}}c_{l_{1}l_{2}l_{3}}^{\mathrm{unfiltered}}$. For the choice of filtering in the simulations presented in Sec. IV, we find that the number of unfiltered coefficients $N_{l,\mathrm{uf}}$ is related to the number of total coefficients $N_{l}$ in each direction as $N_{l,\mathrm{uf}}=N_{l}-1$. Note that the coefficients are always computed using the DG grid points. This is straightforward in an element using the DG scheme, as we have the values of the variable we are checking already at the DG grid points. However, if we want to check the criterion in a FV/FD element, we first interpolate onto the DG grid points using a twelfth-order Lagrange interpolation scheme, and then compute the expansion coefficients on this reconstructed DG grid. Lastly, $\alpha_{N}^{\mathrm{evo}}$ is a relaxation parameter. For our simulations, we set $\alpha_{N}^{\mathrm{evo}}=8$ in the DG elements and $\alpha_{N}^{\mathrm{evo}}=9$ in the FV/FD elements, except for the blast wave propagation tests in Sec. IV.2. We relax our parameters in Sec. IV.2, by using $4$ for DG elements and $5$ for FV/FD elements. The higher values in FV/FD elements is because we want to be more cautious when switching from FV/FD to DG as DG will not work well in a troubled element. From our experience, this Persson trouble indicator plays an especially crucial role in the accuracy of the simulation.
These criteria are checked in order, and if any one of them indicates trouble at any point in an element, the element is marked as troubled by the TEI and no further check is performed.
iii. If an element is marked troubled by the TEI, then if we are using DG, we switch to FV/FD and redo the full RK step (not just the most recent substep). Conversely, if we are using FV/FD then we keep doing FV/FD for the next RK step. On the other hand, if the element is not marked troubled, then if we are using DG, we try doing DG for the next step. However, if we are using FV/FD, then we switch to DG only after 11 consecutive time steps with no trouble to be safe. In the DG elements, we use half the number of points as in the FV/FD elements, as DG can generate higher accuracy results at low computational cost.
@@ IV Numerical results
Table 1: Specifics of the different single neutron stars we simulate. The first column states the name and the abbreviation we assign. The second column states the initial central density. The third and fourth column state the rest mass and the initial star radius (in Schwarzschild coordinates). The fifth column states the strength of the initial perturbation (either in pressure $P$ or density $\rho_{0}$) and the sixth column states the speed imparted to the CoM. The last column states the choices of evolution scheme we use to simulate the case. In each case, we use the ideal EoS of Eq. (25) with $\kappa=100$ and $n=1$.
Name $\rho_{0,\mathrm{C}}\cdot 10^{3}M_{\odot}^{2}$ $m_{0}/M_{\odot}$ $R_{\mathrm{Sch}}/M_{\odot}$ $\lambda$ $v_{\mathrm{CoM}}$ Scheme(s)
@@
In this paper we present simulations that we have performed to ensure the robustness and stability of our new methods. The results of these simulations are presented here through a number of plots; the data related to them can be found in [119]. We start with a scalar wave test, as this helps us test our new DG+FV/FD scheme purely by itself, without the numerical complexities arising from matter fields in neutron star systems. Having checked proper convergence for this simple case, we move on to evolving initial data with special relativistic Riemann discontinuities in the initial data to test hydrodynamic shock propagation. Next, we simulate single neutron stars. We simulate four different cases of single neutron star initial data that are outlined in Table 1, while the resolutions used for each of the initial setup are described in Table 2.
@@
Table 2: Initial resolution, number and type of elements across the star for different single neutron star cases we simulate. The first column is the name of the case as per Table 1 and the second column is the corresponding section where the case is discussed. $L$ in the third column states the dimension of the central cube. The fourth column states the evolution scheme or schemes used. The fifth and sixth columns state the number of points in each direction in DG and FV/FD elements respectively, as applicable. The seventh column states center velocity $v_{\mathrm{CoM}}$. The eighth column specifies the highest level of h-refinement in the refinement region centered on the star. The last column states the initial total number of elements across the star, while specifying how many are of each type. When counting these elements, we also count those which partially contain the star at its surface. In the DG+FV/FD cases, the FV/FD elements are on the star surface, while DG elements are in the center. The elements for the $v_{\mathrm{CoM}}=0.5$ case are counted parallel ($\parallel$) to $\vec{v}_{\mathrm{CoM}}$ direction and perpendicular ($\perp$) to $\vec{v}_{\mathrm{CoM}}$ direction.
Name Sec. $L$ Scheme nn(DG) nn(FV/FD) $v_{\mathrm{CoM}}$ ll(max) Star elements
(S) IV.3.1 $32$ DG 55 - 0 44 $10$ DG
55 $18$ DG
66 $34$ DG
IV.3.2 $52$ DG+FV/FD 88 $16$ 0 55 88 DG ++ 44 FV/FD
(P) IV.4 $32$ DG 55 - 0 44 $10$ DG
55 $18$ DG
66 $34$ DG
(U) IV.5.1 $16$ FV/FD - 1212,1818,2828,$42$ 0 33 55 FV/FD
IV.5.2 $16$ FV/FD - $16$ 0 33 55 FV/FD
44 99 FV/FD
55 $17$ FV/FD
IV.5.3 $16$ DG+ FV/FD 88 $16$ 0 33 55 FV/FD
44 99 FV/FD
(B) IV.6 $52$ DG+FV/FD 88 $16$ $0.5$ 55 55 DG ++ 22 FV/FD $\parallel\vec{v}_{\mathrm{CoM}}$,
55 DG ++ 44 FV/FD $\perp\vec{v}_{\mathrm{CoM}}$
$0.1$ 44 11 DG ++ 44 FV/FD
55 55 DG ++ 44 FV/FD
@@ 55 DG ++ 44 FV/FD $\perp\vec{v}_{\mathrm{CoM}}$
Each of these cases are discussed in detail in the sections below. Out of these, (S) tests for stability and robustness, while the rest test the ability of the program to handle various types of dynamic scenarios. In order to study the behavior of the numerical results with increasing resolution, we use both p-refinement and h-refinement. P-refining an element simply involves increasing the number of points in the element in each direction as necessary. In case of h-refinement of an element, we take the element and split it in two in each direction. This results in eight new elements replacing the old element. Each of these eight new elements contains the same arrangement of points as the parent element. For DG elements, p-refinement increases the basis polynomial order, which would cause an exponential dropoff in the error in the results in the case of smooth fields, but can also prove problematic when nonsmooth fields are present in an element, as these can give rise to sharp noisy features in the higher basis modes. In this case, for DG elements, it is more convenient to increase resolution using h-refinemet instead. For the case of pure FV/FD evolutions on the other hand, we can employ both p-refinement and h-refinement as necessary.
@@
To the best of our knowledge, this is the first time DG methods have been used for boosted neutron stars (B) and for unstable neutron stars (U) (that migrate to a stable state) in fully 3D simulations. While [87, 65, 97] evolve single TOV stars with a hybrid DG-FD scheme, they study the initially static configuration and a rotating configuration. Note, however, that [97] also has successfully evolved a binary inspiral and merger in full 3D without any symmetry assumptions, which is a case we have not studied yet with Nmesh. Although simulations of the initially unstable configuration are performed with DG methods in [83], a 1D spherically symmetric implementation is used. This case setup is also studied in [115], where simulations are performed using spherically symmetric and also axisymmetric cartoon methods [120, 121]. As such, none of the prior simulations are in full 3D. Additionally, none of them evolve boosted neutron stars.
@@ IV.1 Tests with a smooth scalar field
In order to keep message passing interface (MPI) communication at a minimum, our particular FV method does not use any extra ghost points (unlike more traditional finite volume codes). Communication has to happen only for points on the element surfaces as in the DG method. This is crucial for good strong scaling. On the other hand, our FV method becomes less accurate near element boundaries, because we need to drop to a lower order FV scheme there, due to the lack of ghost points. We have tested the method by evolving a smooth scalar field, using the same conservative evolution equations as described in Sec. 4.1 of [76]. We choose the mesh to be a 2-dimensional square that is equally divided into nine smaller square patches. In the middle patch we use FV, while the surrounding eight patches use DG. We then uniformly h-refine each patch, so that each of the nine patches is covered by an equal number of elements. In each element we use only four points (in each of the two directions), which is a challenging test for our FV method that works best if the number of points in the interior is larger than the number of element boundary points, because our method falls back to lower accuracy at element boundaries as explained in Sec. III.2.
@@
In Fig. 3 we show how the numerical error converges to zero with resolution. Our method is consistent with second order convergence. While second-order convergence may not sound too impressive, we note that in shocked regions we cannot expect more than first-order convergence in any case. Thus, our method is adequate when used in elements with shocks. Our plan is then to use this FV method only in the elements where the matter is not smooth, while we plan to use the DG method everywhere else. In this way we should get the benefits of both FV and DG as suggested in [87, 65].